can you explain why B is the correct answer and show the work used to get to that answerPlanet X s only moon travels in an elliptical orbit with eccentricity 0.64 and semi-major axis 5x 107 m. It takes 25 hr to co
Answer #1

The time period is given by

T = 2\pi \sqrt{\frac{a^3}{GM}}

Where T is the time period, a is the semi major axis, G is the gravitational constant M is the mass.

\Rightarrow M = 4\pi^2 \frac{a^3}{GT^2} = \frac{4\pi^2 *(5*10^7)^3}{6.67*10^{-11}*(25*3600)^2} = 9.134*10^{24} kg